OpenStudy (anonymous):

the world population is projected to be 6.6 billion in 2040. at the time, there are expected to be approx 2 billion more people in asia than in africa. the population for the rest of the world will be approx 0.1 billion more than half the population of asia. find the projected populations of asia, africa and the rest of the world in 2040.

OpenStudy (anonymous):

for simplicity sake i'm going to take away the 'billion term' we can add it back at the end let asia = x let africa = y let rest of world = z so first off x + y + z = 6.6 since the difference between x and y is 2 x-y = 2 and the last equation (1/2)x+(1/10) = z

OpenStudy (anonymous):

you can use substitution here , or matrix operators (if you know how) to solve the variables x y and z. once you get numbers for x y and z tag the 'billion' behind them and you are left with your final answer! :)

OpenStudy (anonymous):

ok ty matt

OpenStudy (anonymous):

any questions? is that clear?

OpenStudy (anonymous):

yes its clear